OpenStudy (anonymous):

find the geometric mean of 3 and 15 if necessary give the answer in simplest radical form . please show work thank you

OpenStudy (anonymous):

@judith, same as last one take the square root of the product

OpenStudy (anonymous):

rad(3*15)= 3rad5

OpenStudy (anonymous):

only part that might not be trivial is to write in simplest radical form, look for perfect squares inside the radical

OpenStudy (anonymous):

\[\sqrt{3\times 15}=\sqrt{45}=\sqrt{9\times 5}=\sqrt{9}\sqrt{5}=3\sqrt{5}\]